Can you help me with this task?
3x+4y = 5
x-5y = 7
1
Expert's answer
2010-05-18T04:44:12-0400
Here is an answer for your algebra question:
x = 7+5y
3(7+5y)+4y = 5
21+19y = 5
19y = -16
y = -16/19
x = 7-80/19 = 53/19